suppose real GDP is growing at 4 percent, the money supply is growing at 11 percent, the velocity of money is constant and the real interest rate is 6 percent.
a. what is the current inflation rate and nominal interest rate?
b.if the money supply growth rate increases to 15 percent, how will your answers in part (a) change?
1
Expert's answer
2012-09-06T10:05:55-0400
(i) M/P=Y/V
Y - growing by 3% so coeficient = 1.03
M - growing by 10% so coeficient = 1.10
V - constant =1
1.1/P=1.03/1
P=1.1/1.03≈1.068≈+6.8%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.05=(1+i)/(1+0.068)
i=(1.05*1.068)-1
i≈0.1214≈12.14%
______ ______ ______ ______

(ii)
M/P=Y/V
Y - growing by 3% so coeficient = 1.03
M - growing by 15% so coeficient = 1.15
V - constant =1
1.15/P=1.03/1
P=1.15/1.03≈1.1165≈+11.65%

r-real interest rate
i-nominal interest rate
π-inflation
(1+r)=(1+i)/(1+π)
1.05=(1+i)/(1+0.1165)
i=(1.05*1.1165)-1
i≈0.1723≈17.23%
______ ______ ______ ______

i¹≈12.14%
i²≈17.23%
Δi=(17.23/12.14)-1≈
≈1.4195-1≈+41.95%

Nominal interest rate changed for ≈+41.95% or ≈+5.09 points.